The following net transaction accounts have been documented by a bank for the computation of its reserve requirements (in millions).

Tuesday
11th
Wednesday
12th
Thursday
13th
Friday
14th
Monday
17th
Tuesday
18th
Wednesday
19th
Thursday
20th
Friday
21th
Monday
24th
Net transaction accounts $ 370 $ 320 $ 350 $ 330 $ 350 $ 370 $ 340 $ 330 $ 320 $ 310


The average daily reserves at the Fed for the 14-day reserve maintenance period have been $30.3 million per day, and the average vault cash for the computation period has been estimated to be $3 million per day.

a. What is the amount of the average daily required reserves to be held by the bank during the maintenance period?
b. Is the bank in compliance with the requirements?
Answer #1

a). Average net transaction accounts over the 14-day period = (370+320+350+330+330+330+350+370+340+330+320+320+320+310)/14 = 335

Note: The reserves on Friday are carried over on Saturday and Sunday.

Reserves requirement = sum of [daily average net transaction accounts*reserve percentage]

= sum of [ (range of daily average net transaction accounts)*reserve percentage]

Daily average net transaction accounts Range Required reserve %age Daily average reserves required
8.5 8.5 0% 0.000
45.8-8.5 37.3 3% 1.119
335-45.8 289.2 10% 28.920
30.039

Minimum reserves requirement is $30.039

Daily average requirement = minimum reserves requirement - average vault cash = 30.039 - 3 = 27.039 million

The bank needs to maintain $27.039 million during the maintenance period.

b). Yes, the bank is in compliance because it needs to maintain $27.039 million average reserves whereas it has average reserves of $30.3 million at the Fed.